Please help me to solve this problem....
Angel is twice as old as Gavin.In three years,Angels will be13years less than twice as old as Gavin.
Find their ages

Answer by ankor@dixie-net.com(22740)

let a = A's present age
let g = G's present age
;
Write an equation for each statement:
:
Angel is twice as old as Gavin.
a = 2g
:
In three years, Angels will be 13 years less than twice as old as Gavin.
a + 3 = 2(g+3) - 13
a + 3 = 2g + 6 - 13
a + 3 = 2g - 7
a = 2g - 7 - 3
a = 2g - 10
Replace a with 2g
2g = 2g - 10
Obviously this is not possible, an undo-able problem